Method of Elimination solving problem
14. Solve each linear system.
a) 0.2x-0.3y=1.3
0.5x+0.2y=2.3


b) 0.1a-0.4b=1.9
0.4a+0.5b=-0.8
Pleaseeeeeee and thanksssssss a lottttt

Answer by checkley77(12844)

a) 0.2x-0.3y=1.3 multiply by 2
0.5x+0.2y=2.3 multiply by 3 & add.
.4x-.6y=2.6
1.5x+.6y=6.9
----------------
1.9x=9.5
x=9.5/1.9
x=5 ans.
.2*5-.3y=1.3
1-.3y=1.3
-.3y=1.3-1
-.3y=.3
y=.3/-.3
y=-1 ans.
Proof:
.5*5+.2*-1=2.3
2.5-.2=2.3
2.3=2.3
-------------------------
b) 0.1a-0.4b=1.9 multiply by -4 & add.
0.4a+0.5b=-0.8
-.4a+1.6b=-7.6
-----------------------
2.1b=-8.4
b=-8.4/2.1
b=-4 ans.
.1a-4*-.4=1.9
.1a+1.6=1.9
.1a=1.9-1.6
.1a=.3
a=.3/.1
a=3 ans.
proof:
.4*3+.5*-4=-.8
1.2-2=-.8
.8=.8